OpenStudy (anonymous):

Candice and calra leave the same place at the same time and rode their bikes in the same direction. Candice bikes at an averae speed that was 3/4 Carla speed. After 2 hers they were 28 miles apart. What are their average speed>?

OpenStudy (anonymous):

v1 = (3/4) v2 distance = (velocity )(time) s1 = v1 t s2 = v2 t (s2-s1) = 28 = (v2 - v1) (2) 28 = (v2 - 0.75 v2)(2) 14 = .25 v2 v2 = 56 mph v1 = 42 mph check: v2-v1 = 14 mph (v2-v1)(2hr) = 28 miles